# InfoTok: Adaptive Discrete Video Tokenizer via Information-Theoretic Compression

### v1

@@ 6 Discussions & Limitations
Generalizability of InfoTok.While this work has concentrated on the video domain, the core principles of InfoTok are not intrinsically limited to visual sequences. Our framework is founded on information-theoretic compression and utilizes an ELBO-based router to estimate information complexity, a methodology that is broadly applicable to any data modality compatible with a VAE-style reconstruction objective. The fundamental challenge of variable information density exists in many other domains. For instance, audio signals often contain significant temporal redundancy (e.g., periods of silence, sustained background noise), where an adaptive approach could yield substantial compression gains by allocating tokens based on acoustic complexity (Zhang et al., 2023; Liu et al., 2025). Likewise, 3D data, such as point clouds or meshes, often features sparse data and non-uniform detail, with large, simple surfaces requiring less representational capacity than areas of high geometric complexity (Chen et al., 2025; Lu et al., 2025). Therefore, applying InfoTok to develop adaptive tokenizers for audio, 3D, or other structured data presents a promising direction for future research.
@@ Appendix B Experimental Details
The complete hyperparameter settings of the InfoTok model architecture are provided in Table 5. The adaptive compressor and de-compressor are implemented using a standard Vision Transformer (ViT) equipped with 2D Rotary Position Embeddings (RoPE)Heo et al. (2024) to handle videos with different resolution. Features obtained from the Cosmos encoder are processed by an eight-layer ViT encoder and subsequently compressed using ELBO-guided length selection and likelihood-based token selection, as described in Sections 2 and 3.2. The compressed discrete tokens are decoded back into continuous features via an eight-layer ViT decoder and passed through the Cosmos Tokenizer’s decoder to reconstruct the original video. Despite its effectiveness, the adaptive compressor/de-compressor module constitutes only 14.6% of the total parameters of InfoTok, with the base tokenizer accounting for the remaining 85.4%. This highlights the efficiency and potential of our adaptive tokenization framework.
@@
Positional embedding 2D RoPE Heo et al. (2024) -
@@
Model type FSQ Mentzer et al. (2023b) -
@@
Based Tokenizer: Cosmos Agarwal et al. (2025)
@@
The training augmentation follows Cosmos Tokenizer Agarwal et al. (2025). We employ the AdamW Loshchilov & Hutter (2017) optimizer with an initial learning rate 1e41e-4 with cosine decay to 1e51e-5 for $1e5$ steps. The training regime features a batch size of 1 with video clips of 33 frames over $2e5$ steps. All models presented in this paper are trained with 32 H100 GPUs (4 nodes), and our setting will take 4 days to train a model. Notice that each GPU can consume data with different resolution simultaneously.
@@ B.3 Evaluation Details
TokenBench Agarwal et al. (2025) is a curated benchmark designed to standardize the evaluation of video tokenizers. It comprises 500 high-resolution, long-duration videos sampled from diverse datasets, including BDD100K Yu et al. (2020), EgoExo-4D Grauman et al. (2024), BridgeData V2 Walke et al. (2023), and Panda-70M Chen et al. (2024). These videos encompass a wide range of complex scenarios such as autonomous driving, egocentric activities, robotic manipulation, and web videos, providing a comprehensive evaluation suite for video tokenization methods. For the DAVIS dataset Caelles et al. (2019), we utilize the Test-Dev 2019 split, which consists of 30 video sequences featuring various human activities. As mentioned in Section 4.1, all videos are resized and cropped into 256px to ensure fair comparison with baselines Yan et al. (2024) only supporting square video data tokenization.
@@ Appendix C Additional Experiment Results
Appendix D The Use of Large Language Models
We use Large Language Model to aid and polish writing, e.g., grammar checking and fluency improvement.

### v2

@@ 6 Discussions & Limitations
Generalizability of InfoTok.While this work has concentrated on the video domain, the core principles of InfoTok are not intrinsically limited to visual sequences. Our framework is founded on information-theoretic compression and utilizes an ELBO-based router to estimate information complexity, a methodology that is broadly applicable to any data modality compatible with a VAE-style reconstruction objective. The fundamental challenge of variable information density exists in many other domains. For instance, audio signals often contain significant temporal redundancy (e.g., periods of silence, sustained background noise), where an adaptive approach could yield substantial compression gains by allocating tokens based on acoustic complexity (zhang2023speechtokenizer; liu2025unitok). Likewise, 3D data, such as point clouds or meshes, often features sparse data and non-uniform detail, with large, simple surfaces requiring less representational capacity than areas of high geometric complexity (chen2025meshanything; lu2025uni). Therefore, applying InfoTok to develop adaptive tokenizers for audio, 3D, or other structured data presents a promising direction for future research.
@@ Appendix B Experimental Details
The complete hyperparameter settings of the InfoTok model architecture are provided in Table 4. The adaptive compressor and de-compressor are implemented using a standard Vision Transformer (ViT) equipped with 2D Rotary Position Embeddings (RoPE)heo2024rotary to handle videos with different resolution. Features obtained from the Cosmos encoder are processed by an eight-layer ViT encoder and subsequently compressed using ELBO-guided length selection and likelihood-based token selection, as described in Sections 2 and 3.2. The compressed discrete tokens are decoded back into continuous features via an eight-layer ViT decoder and passed through the Cosmos Tokenizer’s decoder to reconstruct the original video. Despite its effectiveness, the adaptive compressor/de-compressor module constitutes only 14.6% of the total parameters of InfoTok, with the base tokenizer accounting for the remaining 85.4%. This highlights the efficiency and potential of our adaptive tokenization framework.
@@
Positional embedding 2D RoPE heo2024rotary -
@@
Model type FSQ mentzer2023fsq -
@@
Based Tokenizer: Cosmos agarwal2025cosmos
@@
The training augmentation follows Cosmos Tokenizer agarwal2025cosmos. We employ the AdamW loshchilov2017adamw optimizer with an initial learning rate 1e41e-4 with cosine decay to 1e51e-5 for $1e5$ steps. The training regime features a batch size of 1 with video clips of 33 frames over $2e5$ steps. All models presented in this paper are trained with 32 H100 GPUs (4 nodes), and our setting will take 4 days to train a model. Notice that each GPU can consume data with different resolution simultaneously.
@@ B.3 Evaluation Details
TokenBench agarwal2025cosmos is a curated benchmark designed to standardize the evaluation of video tokenizers. It comprises 500 high-resolution, long-duration videos sampled from diverse datasets, including BDD100K yu2020bdd100k, EgoExo-4D grauman2024ego, BridgeData V2 walke2023bridgedata, and Panda-70M chen2024panda. These videos encompass a wide range of complex scenarios such as autonomous driving, egocentric activities, robotic manipulation, and web videos, providing a comprehensive evaluation suite for video tokenization methods. For the DAVIS dataset sergi2019davis, we utilize the Test-Dev 2019 split, which consists of 30 video sequences featuring various human activities. As mentioned in Section 4.1, all videos are resized and cropped into 256px to ensure fair comparison with baselines yan2024elastictok only supporting square video data tokenization.
@@ Appendix C Additional Experiment Results
Appendix D Use of Large Language Models
We use large language models to polish our writing, e.g., to check grammar and improve paper fluency.